# The Limits of Assumption-free Tests for Algorithm Performance

### v1

@@ 2.2 Additional related literature
Many statistical tests and heuristic methods have been proposed for algorithm evaluation and comparison. Given a large dataset, using a holdout set, also known as sample splitting, is often recommended to evaluate or compare algorithms in practice (Hinton et al.,, 1995; Neal,, 1998) (note that using a holdout set to evaluate a fitted model ${\widehat{f}}_{n}$ is directly addressing the question EvaluateModel, rather than EvaluateAlg). With a limited sample size, resampling methods are more popular, including methods such as cross-validation, and its variants such as $5\times 2$-fold cross-validation (CV) (Dietterich,, 1998), (corrected) repeated train-validation (Nadeau and Bengio,, 2003), and nested cross-validation (Varma and Simon,, 2006; Raschka,, 2018). As discussed earlier, many works in the literature have emphasized the qualitative difference between EvaluateAlg and EvaluateModel (e.g., Dietterich, (1998); Hastie et al., (2009)). Which of these questions is more useful to answer, may depend on the particular application; for example, Trippe et al., (2023) argued that EvaluateModel may be a more meaningful question in certain settings. However, in settings where EvaluateAlg is the question of interest, how to perform inference on this target is less well understood. Nadeau and Bengio, (2003) identified that ignoring the variability due to the randomness of the training set can severely undermine the validity of inference in EvaluateAlg. Classical results on CV have suggested that CV can consistently determine which of two algorithms is better, but these results are generally asymptotic and require some restrictive conditions on the data distribution (Shao,, 1993; Yang,, 2007). See also Hastie et al., (2009); Arlot and Celisse, (2010) for surveys on model selection via CV. Based on the asymptotic normality of CV error, the recent work Bayle et al., (2020) discussed how to compare the prediction error of averaged models of two algorithms (in the same spirit as EvaluateModel) in the CV procedure. Understanding the estimand of CV is also an active area of research (Bates et al.,, 2023). There are also many other methods available for estimating the prediction error of an algorithm other than CV; see, e.g., Rosset and Tibshirani, (2019) and references therein for work in a nonparametric regression setting.
Our work also contributes to an important line of work on impossibility or hardness results in the statistics, especially in the distribution-free inference setting (Bahadur and Savage,, 1956; Barber et al., 2021a, ; Shah and Peters,, 2020; Barber,, 2020; Medarametla and Candès,, 2021). The closest work to ours is Kim and Barber, (2023), which establishes a hardness result for the problem of constructing an assumption-free test of algorithmic stability.
@@ 3 Theoretical results for algorithm evaluation
In this section, we present our first main result, establishing the hardness of answering the question EvaluateAlg—or more concretely, of performing the hypothesis test (1)—via a black-box procedure. With the aim of avoiding placing any assumptions on the algorithm $\mathcal{A}$ and on the distribution $P$, we would like to provide an answer to EvaluateAlg that is uniformly valid over all possible $\mathcal{A}$ and $P$. Specifically, fixing a error tolerance level $\alpha$, a sample size $n$, an available dataset $\mathcal{D}_{N}=\{(X_{i},Y_{i})\}_{i\in[N]}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and a risk threshold $\tau>0$ for our hypothesis test, we will consider tests ${\widehat{T}}={\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$ that satisfy
@@
where the probability is taken with respect to the data set $\mathcal{D}_{N}$ (as well as any randomization in the test ${\widehat{T}}$). From this point on, we treat the allowed error level $\alpha\in(0,1)$, the risk threshold $\tau>0$, and sample sizes $n,N\geq 1$ as fixed; our results will give finite-sample (non-asymptotic) limits on power for this testing problem. We will see that power for any black-box test is limited unless the available sample size $N$ is much larger than $n$.
3.1 Limits of black-box tests for unbounded loss
First, we consider the setting of an unbounded loss function:
$\sup_{{\widehat{y}}\in\widehat{\mathcal{Y}},y\in\mathcal{Y}}\ell({\widehat{y}},y)=\infty.$
We will see that in this case, no black-box test can have better-than-random power for inference on the risk of any algorithm $\mathcal{A}$.
Theorem 1.
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ is unbounded, i.e., $\sup_{{\widehat{y}},y}\ell({\widehat{y}},y)=\infty$. Let ${\widehat{T}}$ be a black-box test (as in Definition 1), and assume that ${\widehat{T}}$ satisfies the assumption-free validity condition (5). Then the power of ${\widehat{T}}$ to detect low risk is bounded as follows: for any $\mathcal{A}$ and any $P$ with $R_{P,n}(\mathcal{A})<\tau$,
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\leq\alpha.$
In other words, any black-box test is completely powerless for any algorithm $\mathcal{A}$ and distribution $P$—the probability of certifying low risk for $(\mathcal{A},P,n)$ is no better than random. From this point on, then, in this paper we will consider the case of a bounded loss.
Remark 1 (The infinite cardinality assumption).
In this theorem (and in all our results below), we assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, i.e., the data points lie in a space $\mathcal{X}\times\mathcal{Y}$ that has infinite cardinality. The reason for this assumption is that, for finite $\mathcal{X}\times\mathcal{Y}$, since the definition of a black-box test does not place any computational constraints on the number of calls to $\mathcal{A}$, we can perform an exhaustive search to fully characterize the behavior of $\mathcal{A}$. Therefore, the problem of testing the risk of $\mathcal{A}$ with no computational constraints is meaningful only for an infinite-cardinality space. On the other hand, in practice computational limits are a natural component of any testing problem, and therefore inference on risk for a large but finite space $\mathcal{X}$ likely will have power limited by some function of the computational budget. We leave this more complex question for future work. (See also Appendix A.1 for further discussion.)
@@ 3.2 Limits of black-box tests for bounded loss
the largest possible risk of any function $f$ for data drawn from $P$. We are primarily interested in testing the hypothesis (1) for a threshold $\tau\leq R_{P}^{\max}$ (because if we were instead considering $\tau>R_{P}^{\max}$, this would be a fundamentally simpler setting—any algorithm will have risk $<\tau$ over distribution $P$, so the hypothesis test would depend only on $P$, not on the algorithm $\mathcal{A}$).
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ takes values in $[0,B]$. Let ${\widehat{T}}$ be a black-box test (as in Definition 1), and assume that ${\widehat{T}}$ satisfies the assumption-free validity condition (5). Let $\tilde{\tau}=\tau(1+\frac{\alpha^{-1}-1}{N})$.
Then the power of ${\widehat{T}}$ to detect low risk is bounded as follows: for any $\mathcal{A}$ and any $P$ with $R_{P,n}(\mathcal{A})<\tau$,
@@
In particular, as expected, if the risk $R_{P,n}(\mathcal{A})$ is only slightly lower than $\tau$, then the power is only slightly higher than $\alpha$ (i.e., no better than random). But, even if the risk is low (e.g., $R_{P,n}(\mathcal{A})=0$ at the extreme), if the available data is limited (e.g., $N\propto n$), then we see that the power can only be a constant multiple of $\alpha$. If the allowed error level $\alpha$ is small, we therefore cannot have a powerful test for any algorithm $\mathcal{A}$.
Conversely, given a data set size $N$, if we wish to have high power for identifying whether an algorithm $\mathcal{A}$ has low risk $R_{P,n}(\mathcal{A})$, we need to choose a small sample size, $n\ll N$—but for small $n$, the algorithm $\mathcal{A}$ will likely return fitted models that are less accurate, i.e., the risk $R_{P,n}(\mathcal{A})$ itself will likely be higher. (This type of tradeoff has been described by Rinaldo et al., (2019) as the inference–prediction tradeoff, in a related setting.)
@@
The results of Theorem 2 bound the power of any black-box test ${\widehat{T}}$ when applied to any algorithm $\mathcal{A}$, including algorithms that may be randomized (i.e., $\mathcal{A}(\mathcal{D};\xi)$ has nontrivial dependence on the random seed $\xi$). This is a stronger conclusion than if we only bounded power for deterministic algorithms, but the assumption is stronger as well: the validity condition (5) requires ${\widehat{T}}$ to have a bounded error rate over all $\mathcal{A}$ (including algorithms that are randomized). We remark that the deterministic version of the result also holds: that is, if we require validity (5) only with respect to deterministic $\mathcal{A}$ (algorithms for which $\mathcal{A}(\mathcal{D};\xi)$ does not have any dependence on $\xi$), then the bound on power given in Theorem 2 also holds for all deterministic $\mathcal{A}$ (in fact, the same construction in the proof can be used for both versions of the result).
@@ 3.3 A matching bound
and evaluate the loss on a single test point,
@@
Comparing this power calculation to the result of Theorem 2, we can see that in a setting where $n,N$ are large and $R_{P}^{\max}\approx 1$, the power of the simple test ${\widehat{T}}_{\textnormal{Binom}}$ is nearly as high as the optimal power that can be achieved by any black-box test. This may seem counterintuitive: how can this simple test be (nearly) optimal, given that it uses the data very inefficiently—without any more sophisticated strategies such as cross-validation, resampling, synthetic or perturbed data, etc? Nonetheless, comparing these theorems reveals that the simple Binomial test, based only on sample splitting, is nearly optimal because of the fundamental hardness of the problem.
A concrete example.
@@ 4 The role of algorithmic stability
Theorem 2 suggests that to alleviate the hardness result for algorithm evaluation, we must consider placing assumptions on either the data distribution $P$ or the algorithm $\mathcal{A}$, since any universally valid test ${\widehat{T}}$ must necessarily have low power.
A large body of recent works suggest that in many contexts, algorithm stability helps CV-based methods achieve distribution-free guarantees (Kumar et al.,, 2013; Barber et al., 2021b, ; Austern and Zhou,, 2020; Bayle et al.,, 2020). This suggests that this might also be the case for the problem of addressing EvaluateAlg: is estimating the risk $R_{P,n}(\mathcal{A})$ achievable with a stability assumption, and no other conditions? In this section, we will explore this possibility. Perhaps surprisingly, we will see that the hardness result of Theorem 2 remains the same even with a stability assumption, except in an extremely high-stability regime.

### v4

@@ 2.2 Additional related literature
Many statistical tests and heuristic methods have been proposed for algorithm evaluation and comparison. Given a large dataset, using a holdout set, also known as sample splitting, is often recommended to evaluate or compare algorithms in practice (Hinton et al.,, 1995; Neal,, 1998) (note that using a holdout set to evaluate a fitted model ${\widehat{f}}_{n}$ is directly addressing the question EvaluateModel, rather than EvaluateAlg). With a limited sample size, resampling methods are more popular, including methods such as cross-validation, and its variants such as $5\times 2$-fold cross-validation (CV) (Dietterich,, 1998), (corrected) repeated train-validation (Nadeau and Bengio,, 2003), and nested cross-validation (Varma and Simon,, 2006; Raschka,, 2018). As discussed earlier, many works in the literature have emphasized the qualitative difference between EvaluateAlg and EvaluateModel (e.g., Dietterich, (1998); Hastie et al., (2009)). Which of these questions is more useful to answer, may depend on the particular application; for example, Trippe et al., (2023) argued that EvaluateModel may be a more meaningful question in certain settings. However, in settings where EvaluateAlg is the question of interest, how to perform inference on this target is less well understood. Nadeau and Bengio, (2003) identified that ignoring the variability due to the randomness of the training set can severely undermine the validity of inference in EvaluateAlg. Classical results on CV have suggested that CV can consistently determine which of two algorithms is better, but these results are generally asymptotic and require some restrictive conditions on the data distribution (Shao,, 1993; Yang,, 2007). See also Hastie et al., (2009); Arlot and Celisse, (2010) for surveys on model selection via CV. Based on the asymptotic normality of CV error, the recent work Bayle et al., (2020) discussed how to compare the prediction error of averaged models of two algorithms (in the same spirit as EvaluateModel) in the CV procedure. Similar results have been also extended to compare the algorithm’s risk (Austern and Zhou,, 2020; Li,, 2023). Understanding the estimand of CV and effectiveness of CV in evaluating a fitted model are also active areas of research (Bates et al.,, 2023; Iyengar et al.,, 2025). There are also many other methods available for estimating the prediction error of an algorithm other than CV; see, e.g., Rosset and Tibshirani, (2019) and references therein for work in a nonparametric regression setting.
Our work also contributes to an important line of work on impossibility or hardness results in statistics, especially in the distribution-free inference setting (Bahadur and Savage,, 1956; Barber et al., 2021a, ; Shah and Peters,, 2020; Barber,, 2020; Medarametla and Candès,, 2021; Kim et al.,, 2022). The closest work to ours is Kim and Barber, (2023), which establishes a hardness result for the problem of constructing an assumption-free test of algorithmic stability.
@@ 3 Theoretical results for algorithm evaluation
In this section, we present our first main result, establishing the hardness of answering the question EvaluateAlg—or more concretely, of performing the hypothesis test (1)—via a black-box procedure. With the aim of avoiding placing any assumptions on the algorithm $\mathcal{A}$ and on the distribution $P$, we would like to provide an answer to EvaluateAlg that is uniformly valid over all possible $\mathcal{A}$ and $P$. Specifically, fixing an error tolerance level $\alpha$, an evaluation sample size $n$, an available dataset $\mathcal{D}_{N}=\{(X_{i},Y_{i})\}_{i\in[N]}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$ of size $N$, and a risk threshold $\tau>0$ for our hypothesis test, we will consider tests ${\widehat{T}}={\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$ that satisfy
@@
where the probability is taken with respect to the data set $\mathcal{D}_{N}$ (as well as any randomization in the test $^{\widehat{T}}$). From this point on, we treat the allowed error level $\alpha\in(0,1)$, the risk threshold $\tau>0$, and sample sizes $n,N\geq 1$ as fixed; our results will give finite-sample (non-asymptotic) limits on power for this testing problem. We will see that power for any valid black-box test for solving (1) is limited unless the available sample size $N$ is much larger than $n$.
@@ 3.1 Limits of black-box tests for bounded loss
the largest possible risk of any function $f$ for data drawn from $P$. We are primarily interested in testing the hypothesis (1) for a threshold $\tau\leq R_{P}^{\max}$ (because if we were instead considering $\tau>R_{P}^{\max}$, this would be a fundamentally simpler setting—any algorithm will have risk $\tau$ over distribution $P$, so the hypothesis test would depend only on $P$, not on the algorithm $\mathcal{A}$).
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ takes values in $[0,B]$. Let $^{\widehat{T}}$ be a black-box test (as in Definition 1), and assume that $^{\widehat{T}}$ satisfies the assumption-free validity condition (6). Let $\tilde{\tau}=\tau(1+\frac{\alpha^{-1}-1}{N})$.
Then the power of $^{\widehat{T}}$ to detect low risk is bounded as follows: for any $\mathcal{A}$ and any $P$ with $R_{P,n}(\mathcal{A})<\tau$,
@@
In particular, as expected, if the risk $R_{P,n}(\mathcal{A})$ is only slightly lower than $\tau$, then the power is only slightly higher than $\alpha$ (i.e., no better than random). But, even if the risk is low (e.g., $R_{P,n}(\mathcal{A})=0$ at the extreme), if the available data is limited (e.g., $N$ is close to the evaluation sample size $n$), then we see that the power can only be a constant multiple of $\alpha$. If the allowed error level $\alpha$ is small, we therefore cannot have a powerful test for any alternative. On the other hand, under the same setting as in Theorem 1, we should clarify that EvaluateModel is relatively easy, as a test leveraging the holdout data can solve the problem. For instance if $N=n+m$, then we have a holdout set of $m$ many data points with which we can estimate $R_{P}({\widehat{f}}_{n})$ up to error $\mathcal{O}_{P}(\frac{1}{\sqrt{m}})$ by concentration—and therefore, testing whether $R_{P}({\widehat{f}}_{n})<\tau$ is straightforward in the black-box setting. We would like to emphasize that Theorem 1 is a hardness result and it is not obvious a priori given many successful examples in distribution-free inference (Vovk et al.,, 2005; Angelopoulos et al., 2024a, ). This suggests that it may be interesting to consider different definitions of assumption-free validity, since the definition of validity given in (6) leads to the above hardness results and consequently appears to be too strong for meaningful inference to be possible.
In addition, the above result has a few other important implications. First, it complements the results in Dietterich, (1998); Nadeau and Bengio, (2003) and provides more quantitative results on the difference between EvaluateAlg and EvaluateModel. Second, it reveals that in the assumption-free setting with a limited number of samples, any claims about the risk of a procedure should be interpreted as a claim about the risk of the fitted model rather than the risk of the algorithm. Third, it reveals a trade-off in our ability to perform inference on EvaluateAlg. This is because given a data set size $N$, if we wish to have high power for identifying whether an algorithm $\mathcal{A}$ has low risk $R_{P,n}(\mathcal{A})$, we need to choose a small sample size, $n\ll N$—but for small $n$, the algorithm $\mathcal{A}$ will likely return fitted models that are less accurate, i.e., the risk $R_{P,n}(\mathcal{A})$ itself will likely be higher. (This type of tradeoff has been described by Rinaldo et al., (2019) as the inference–prediction tradeoff, in a related setting.)
Remark 1.
We note that the bound for the loss function $B$ does not explicitly appear in Theorem 1, but the value of $\tau,R_{P,n}(\mathcal{A})$ and $R_{P}^{\max}$ typically scales with $B$. For example, in common settings, if we multiply the loss function by $10$, then $\tau,R_{P,n}(\mathcal{A})$ and $R_{P}^{\max}$ will also increase by a multiplicative factor $10$. But in that case, the ratio $\frac{\tilde{\tau}-R_{P,n}(\mathcal{A})}{R_{P}^{\max}-\tilde{\tau}}$ in Theorem 1 would not change. In that sense, the power upper bound we get is scale-invariant.
@@
The results of Theorem 1 bound the power of any black-box test $^{\widehat{T}}$ when applied to any algorithm $\mathcal{A}$, including algorithms that may be randomized (i.e., $\mathcal{A}(\mathcal{D};\xi)$ has nontrivial dependence on the random seed $\xi$). This is a stronger conclusion than if we only bounded power for deterministic algorithms, but the assumption is stronger as well: the validity condition (6) requires $^{\widehat{T}}$ to have a bounded error rate over all $\mathcal{A}$ (including algorithms that are randomized). We remark that the deterministic version of the result also holds: that is, if we require validity (6) only with respect to deterministic $\mathcal{A}$ (algorithms for which $\mathcal{A}(\mathcal{D};\xi)$ does not have any dependence on $\xi$), then the bound on power given in Theorem 1 also holds for all deterministic $\mathcal{A}$ (in fact, the same construction in the proof can be used for both versions of the result).
Remark 3 (The infinite cardinality assumption).
In this theorem (and in all our results below), we assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, i.e., the data points lie in a space $\mathcal{X}\times\mathcal{Y}$ that has infinite cardinality. The reason for this assumption is that, for finite $\mathcal{X}\times\mathcal{Y}$, since the definition of a black-box test does not place any computational constraints on the number of calls to $\mathcal{A}$, we can perform an exhaustive search to fully characterize the behavior of $\mathcal{A}$. Therefore, the problem of testing the risk of $\mathcal{A}$ with no computational constraints is meaningful only for an infinite-cardinality space. On the other hand, in practice computational limits are a natural component of any testing problem, and therefore inference on risk for a large but finite space $\mathcal{X}$ likely will have power limited by some function of the computational budget. We leave this more complex question for future work. (See also Appendix B.1 for further discussion.)
Careful readers may wonder does the result in Theorem 1 hold for testing other notions of algorithm risk. One possible choice is the following PAC-type notion: given any $\delta\in(0,1),\epsilon>0$, consider
$H_{0}:\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)\geq\delta\quad\textnormal{versus}\quad H_{1}:\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)<\delta.$ (7)
Then we can ask what is the best power we can hope for if the test is required to satisfy the following notion of assumption-free validity:
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\leq\alpha\textnormal{ for any $\mathcal{A},P$ such that $\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)\geq\delta$}.$ (8)
It turns out that the hardness result established in Theorem 1 also transfers to the new testing problem (7).
Proposition 1.
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ takes values in $[0,B]$. Let $^{\widehat{T}}$ be a black-box test (as in Definition 1), and assume that $^{\widehat{T}}$ satisfies the assumption-free validity condition (8). Let $\tau=\delta B+\epsilon$ and $\tilde{\tau}=\tau(1+\frac{\alpha^{-1}-1}{N})$.
Then the power of $^{\widehat{T}}$ to detect low risk in the sense of (7) is bounded as follows: for any $\mathcal{A}$ and any $P$ with $\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)<\delta$,
@@ 3.2 A matching bound
and evaluate the loss on the last data point of this batch,
@@
Comparing this power calculation to the result of Theorem 1, we can see that in a setting where $n,N$ are large and $R_{P}^{\max}\approx 1$, which is typically the case for loss functions mentioned above as for each $\widehat{y}$, there is always a $y$ such that $\ell({\widehat{y}},y)=1$, the power of the simple test ${\widehat{T}}_{\textnormal{Binom}}$ is nearly as high as the optimal power that can be achieved by any black-box test. This may seem counterintuitive: how can this simple test be (nearly) optimal, given that it uses the data very inefficiently—without any more sophisticated strategies such as cross-validation, resampling, synthetic or perturbed data, etc? Nonetheless, comparing Theorem 1 and Theorem 2 reveals that the simple Binomial test, based only on sample splitting, is nearly optimal because of the fundamental hardness of the problem.
3.2.1 A concrete example.
@@
3.3 A note for unbounded loss
Curious readers may wonder what would happen if the loss is unbounded, i.e., $\sup_{{\widehat{y}}\in\widehat{\mathcal{Y}},y\in\mathcal{Y}}\ell({\widehat{y}},y)=\infty$. It is not difficult to convenience ourselves that when the loss is unbounded, the problem becomes even harder. For EvaluateAlg, we can show that under the same setting as in Theorem 1, no black-box test can have better-than-random power for inference on the risk of any algorithm $\mathcal{A}$.
Theorem 3.
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ is unbounded, i.e., $\sup_{{\widehat{y}},y}\ell({\widehat{y}},y)=\infty$. Let $^{\widehat{T}}$ be a black-box test (as in Definition 1), and assume that $^{\widehat{T}}$ satisfies the assumption-free validity condition (6). Then the power of $^{\widehat{T}}$ to detect low risk is bounded as follows: for any $\mathcal{A}$ and any $P$ with $R_{P,n}(\mathcal{A})<\tau$,
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\leq\alpha.$
At the same time, when the loss is unbounded, assumption-free inference for EvaluateModel also becomes trickier, as the concentration for the estimator $\frac{1}{|\mathcal{D}_{\textnormal{hold}}|}\sum_{(X_{i},Y_{i})\in\mathcal{D}_{\textnormal{hold}}}\ell({\widehat{f}}_{n}(X_{i}),Y_{i})$ could fail. During the revision, one of the reviewers pointed out that under a slightly stronger assumption on the marginal distribution of $X$, i.e., assuming it is non-atomic, then assumption-free inference for the model risk is also impossible when the loss is unbounded. In this scenario, we also need additional assumptions to infer model risk, see some examples in Bousquet and Elisseeff, (2002); Celisse and Guedj, (2016).
On the other hand, we also find that the bounded loss assumption may not be essential in separating the difficulty of EvaluateAlg and EvaluateModel. In particular, if we slightly change the inference target from mean risk to “median” risk, then the difference of EvaluateAlg and EvaluateModel will manifest again. We refer readers to Appendix A for more details.
@@ 4 The role of algorithmic stability
Theorem 1 suggests that to alleviate the hardness result for algorithm evaluation, we must consider placing assumptions on either the data distribution $P$ or the algorithm $\mathcal{A}$, since any universally valid test $^{\widehat{T}}$ must necessarily have low power.
A large body of recent works suggests that in many contexts, algorithm stability helps CV-based methods achieve distribution-free guarantees (Kumar et al.,, 2013; Barber et al., 2021b, ; Austern and Zhou,, 2020; Bayle et al.,, 2020). This suggests that this might also be the case for the problem of addressing EvaluateAlg: is estimating the risk $R_{P,n}(\mathcal{A})$ achievable with a stability assumption, and no other conditions? In this section, we will explore this possibility. Perhaps surprisingly, we will see that the hardness result of Theorem 1 remains the same, except in an extremely high-stability regime.